OpenStudy (boxman61):

math check please, trying to learn this (√x-4 +3)^2 = x+3-4√x-4

sam (.sam.):

\[(\sqrt{x-4}+3)^{2}??\]

OpenStudy (boxman61):

yes, just trying to figure out this square system

OpenStudy (boxman61):

really clueless...obviously lol

sam (.sam.):

\[(\sqrt{x-4}+3)(\sqrt{x-4}+3)\] FOIL \[(x-4)+3\sqrt{x-4}+3\sqrt{x-4}+9\]

sam (.sam.):

\[x+5+6\sqrt{x-4}\]

OpenStudy (boxman61):

thank you...somehow that never occured to me...

sam (.sam.):

If you dont understand why (x−4). Example \[\huge (\sqrt{x})(\sqrt{x})\] \[\huge (x ^{\frac{1}{2}})(x ^{\frac{1}{2}})\] \[\huge x ^{\frac{1}{2}+\frac{1}{2}}=x\]

OpenStudy (boxman61):

just learned that.... just finished my last problem (that yo and others answered earlier) now trying to learn it...thank you very much